Rockness Recycling refurbishes rundown business students. The process uses a moving belt, which carries each student through the five steps of the process in sequence. The five steps are as follows:

STEP DESCRIPTION TIME REQUIRED
PER STUDENT
1 Unpack and place on belt 0.2 minute
2 Strip off bad habits 0.4 minute
3 Scrub and clean mind 0.6 minute
4 Insert modern methods 0.8 minute
5 Polish and pack 1.0 minutes

One faculty member is assigned to each of these steps. Faculty members work a 40-hour week and rotate jobs each week. Mr. Rockness has been working on a contract from General Eclectic, which requires delivery of 2,500 refurbished students per week. A representative of the human resources department has just called complaining that the company hasn’t been receiving the agreed-upon number of students. A check of finished goods inventory by Mr. Rockness reveals that there is no stock left.

a. Calculate the maximum output for each step.

b. Which step is the bottleneck?

  • step5

  • step1

  • step2

  • step3

  • step4
Answer #1

a.
Ans -
Max Capacity = 40 hrs x 60 mins = 2400 minutes
Step 1 - minute unpack and place on belt
Time required per student = 0.2 min
Max. Output = 2400/0.2 = 12,000 students per week
Step 2 - strip off bad habits
Time required per student = 0.4 min
Max. Output = 2400/0.4 = 6,000 students per week
Step 3 - scrub and clean mind
Time required per student = 0.6 min
Max. Output = 2400/0.6 = 4,000 students per week
Step 4 - insert modern methods
Time required per student =0.8 min
Max. Output = 2400/0.8 = 3000 students per week
Step 5 - polish and pack
Time required per student =1 min
Max. Output = 2400/1 = 2400 students per week
b.
Ans -
Bottleneck: Least capacity in process = STEP 5
Process cannot create > what bottleneck is, so 2400 is max capacity for process as a whole.